# Gravitational lensing by charged black hole with global monopole in strong field limit

### v1

@@ II.3 Typical observables in gravitational lensing
Figure 1: The fundamental geometric configuration underlying the phenomenon of gravitational lensing. $b$ is the impact parameter, $r_{0}$ represents the minimum distance between the light ray and the lens, and $I$ and $L$ denote the positions of the image and the lens. $S$ and $O$ represent the source and observer.
The image formed by photons that encircle a lens $n$ times is referred to as the $n$-th relativistic image. We consider the scenario when the source, lens, and observer are nearly perfectly aligned, which is precisely the condition under which relativistic imaging effects are most prominent. The angular position $\theta_{n}$ and magnification of the $n$-th image are given by Bozza et al. (2001); Bozza (2002)
$\theta_{n}=\frac{b_{m}}{D_{\text{OL}}}(1+e_{n})=\theta_{\infty}(1+e_{n}),\quad e_{n}=e^{\frac{\bar{b}-2n\pi}{\bar{a}}},$ (2.27)
$\mu_{n}=e_{n}\frac{b_{m}^{2}\left(1+e_{n}\right)D_{\text{OS}}}{\bar{a}\beta D_{\text{OL}}^{2}D_{\text{LS}}},$ (2.28)
where $\theta_{\infty}=b_{m}/D_{\text{OL}}$ designates the angular position where relativistic images are formed at the photon sphere.
It is hypothesized that the outermost (first) image in the series of external images stands apart from the others, being uniquely discernible, while the remaining images nearly overlap and are difficult to distinguish. Therefore, we introduce an observable quantity: the angular separation, denoted as $s$, which quantifies the difference between the angular position of the first image and that of the subsequent ones. This quantity is defined as
$s=\theta_{1}-\theta_{\infty}.$ (2.29)
Furthermore, another observable is the ratio of the flux intensity of the primary image to the cumulative flux intensity of all subsequent images (hereafter abbreviated as the flux ratio). It is also known as the relative magnification. This ratio provides valuable insights into the relative brightness of the primary image compared to the collective contribution of the other images. It is given by
@@
The detailed derivations of these observables can be found in Refs. Bozza et al. (2001); Bozza (2002).
III Gravitational lensing of RN black hole with a global monopole
We now investigate the gravitational lensing of an RN black hole with a global monopole. Its line element is given by Barriola and Vilenkin (1989)
$ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}d\phi^{2},$ (3.1)
where
$f(r)=1-8\pi\gamma^{2}-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}.$ (3.2)
$\epsilon=1-8\pi\gamma^{2}=1-\frac{\Delta}{4\pi},$ (3.3)
which ranges from $\epsilon=0$ (corresponding to a maximum deficit solid angle $\Delta=4\pi$) to to $\epsilon=1$ (when no monopole is present, i.e., $\Delta=0$), and the charge parameter $\eta$ as
$\eta=Q^{2}/M^{2}.$ (3.4)
This parameter quantifies the effect of charge, with values ranging from $\eta=0$ (uncharged black hole) to $\eta=1$ (extremal black hole in standard RN spacetime). To ensure that the spacetime does not develop a naked singularity, we impose the constraint $|Q|\leq M/\sqrt{\epsilon}$. For generality, and to avoid naked singularities for any $\epsilon$, we restrict the charge to $|Q|\leq M$, leading to the constraint $\eta\in[0,1]$. In the following sections, we analyze the gravitational lensing characteristics of this black hole spacetime and explore how the interplay between the charge and the global monopole affects lensing observables.
III.1 Physical quantities associated with gravitational lensing
In order to calculate the angular separation $s$ and the flux ratio $\mathcal{R}$ in the strong field limit using Eqs. (2.29) and (2.32), it is necessary to obtain the radius of the photon sphere $r_{m}$, the critical impact parameter $b_{m}$, the coefficients $\bar{a}$ and $\bar{b}$, and the regular integral $I_{R}$. These quantities are computed as follows.
From Eq. (2.5), the radius of the photon sphere $r_{m}$ is given by
$r_{m}\left(\epsilon,\eta\right)=\bar{r}_{m}\left(\epsilon,\eta\right)M,$ (3.5)
where
$\bar{r}_{m}\left(\epsilon,\eta\right)=\frac{3+\sqrt{9-8\epsilon\eta}}{2}.$ (3.6)
Meanwhile, the critical impact parameter $b_{m}$ can be obtained from Eq. (2.6) as
$b_{m}\left(\epsilon,\eta\right)=\frac{\bar{r}_{m}^{2}\left(\epsilon,\eta\right)}{\sqrt{\epsilon\bar{r}_{m}^{2}\left(\epsilon,\eta\right)-2\bar{r}_{m}\left(\epsilon,\eta\right)+\eta}}M.$ (3.7)
Based on Eqs. (2.24) and (2.25), the coefficients of the deflection angle $\bar{a}$ and $\bar{b}$ are derived as
$\bar{a}\left(\epsilon,\eta\right)=\frac{\bar{r}_{m}\left(\epsilon,\eta\right)}{\sqrt{\epsilon\bar{r}_{m}^{2}\left(\epsilon,\eta\right)-2\eta}}$ (3.8)
and
$\bar{b}\left(\epsilon,\eta\right)=\bar{a}\left(\epsilon,\eta\right)\ln\left[\frac{6\bar{r}_{m}\left(\epsilon,\eta\right)-8\eta}{\bar{r}_{m}\left(\epsilon,\eta\right)-\eta}\right]+I_{R}\left(\bar{r}_{m}\right)-\pi,$ (3.9)
respectively, where the regular integral $I_{R}$ can be obtained with Eq. (2.18) as
$\begin{split}I_{R}\left(\bar{r}_{m}\right)&=\int_{0}^{1}{\frac{2\bar{r}_{m}}{z\sqrt{-\eta z^{2}-\left(2\bar{r}_{m}-4\eta\right)z+3\bar{r}_{m}+4\eta}}}-\frac{2\bar{a}}{z}dz\\ &=\bar{a}\ln\frac{4\left(3\bar{r}_{m}-4\eta\right)^{2}}{\left(2\bar{r}_{m}-2\eta+\sqrt{\bar{r}_{m}-\eta}\sqrt{3\bar{r}_{m}-4\eta}\right)^{2}},\end{split}$ (3.10)
with $\bar{r}_{m}$ and $\bar{a}$ representing $\bar{r}_{m}(\epsilon,\eta)$ and $\bar{a}(\epsilon,\eta)$ for brevity. Thus, $\bar{b}$ is given by
$\bar{b}(\epsilon,\eta)=\bar{a}(\epsilon,\eta)\ln\left[\frac{8\left(3\bar{r}_{m}(\epsilon,\eta)-4\eta\right)^{3}}{\bar{r}_{m}^{2}(\epsilon,\eta)\left(\bar{r}_{m}(\epsilon,\eta)-\eta\right)^{2}}\left(2\sqrt{\bar{r}_{m}(\epsilon,\eta)-\eta}-\sqrt{3\bar{r}_{m}(\epsilon,\eta)-4\eta}\right)^{2}\right]-\pi.$ (3.11)
When the global monopole is absent ($\epsilon=1$), the coefficients $\bar{a}$ and $\bar{b}$ recover the results for the RN black hole, as derived in Ref. Tsukamoto (2017). Similarly, in the absence of charge ($\eta=0$), the coefficients $\bar{a}$ and $\bar{b}$ reduce to the expressions obtained for a Schwarzschild black hole with a global monopole, as given in Ref. Cheng and Man (2011). Furthermore, in the absence of both charge and the global monopole ($\eta=0$, $\epsilon=1$), the coefficients $\bar{a}$ and $\bar{b}$ match the results from Ref. Bozza (2002) for a Schwarzschild black hole. These consistency checks further validate our approach, ensuring that the inclusion of both charge and a global monopole provides a unified framework that naturally extends existing results.
III.2 Typical observables in gravitational lensing
@@
By leveraging the coefficients $\bar{a}$ and $\bar{b}$ of the deflection angle, we now derive explicit analytical expressions for the angular separation $s$ and the flux ratio $\mathcal{R}$ using Eqs. (2.29) and (2.32). Previous studies have primarily relied on numerical approaches to analyze these observables in specific scenarios. For instance: Ref. Bozza (2002) conducted numerical computations for an RN black hole, but did not provide analytical formulas for $s$ and $\mathcal{R}$. Ref. Cheng and Man (2011) focused on a Schwarzschild black hole with a global monopole, again without explicit analytical expressions for the observables. Ref. Tsukamoto (2017) derived the coefficients $\bar{a}$ and $\bar{b}$ for the RN black hole but did not explore the associated lensing observables. In this work, we go beyond previous studies by deriving explicit analytical expressions for both $s$ and $\mathcal{R}$ for an RN black hole with a global monopole, facilitating a direct evaluation of how these observables respond to variations in charge ($\eta$) and the global monopole parameter ($\epsilon$). To achieve a comprehensive understanding, we conduct both analytical and numerical analyses, examining both the individual effects of the global monopole and charge on the observables, as well as the combined impact of charge and the global monopole, revealing their interplay in modifying strong-field lensing behavior.
III.2.1 The angular separation
@@
Secondly, we investigate the dependence of the angular separation on the charge for a fixed global monopole parameter. First, when there is charge but no global monopole (i.e., $\epsilon=1$), the angular separation is given by the expression:
@@
where $w$ represents the radius of the photon sphere, $\bar{r}_{m}(1,\eta)$. By substituting the values of charge $Q$ from Ref. Bozza (2002), this reproduces the numerical results for the angular separation in the RN black hole case. Furthermore, as shown in Appendix A, the derivative of the angular separation with respect to $\eta$ is always positive. This indicates that the angular separation monotonically increases with increasing charge. Now, we analyze the behavior of the angular separation $s$ in the presence of both charge and a global monopole. Directly analyzing Eq. (3.12) is challenging due to its complexity. To understand the effects of both charge and the global monopole, we expand the angular separation expression for small and large charge values, respectively. For small charge ($\eta\ll 1$), Eq. (3.12) can be expanded as
@@
The dependence of the angular separation on charge is determined by the coefficient of the $\eta$-dependent term. There exists a critical value of the global monopole parameter, given by $\epsilon_{0}=\left[\frac{51\sqrt{3}-88}{2\pi\left(7\sqrt{3}-12\right)}\right]^{2}\approx 0.1833742$, which separates two behaviors: For $\epsilon>\epsilon_{0}$, coefficient of the $\eta$-dependent term is positive, and the angular separation increases with charge. Conversely, for $\epsilon<\epsilon_{0}$, the coefficient is negative, and the angular separation decreases with charge. For large charge ($\eta\to 1$), the angular separation (Eq. (3.12)) behaves as:
@@
where $H(\epsilon)$ and $F(\epsilon)$ are functions of the global monopole parameter $\epsilon$. Due to the complexity of these expressions, their explicit forms are provided in Appendix B. We find that, similar to the small-charge case, there exists a critical value $\epsilon_{1}\approx 0.1826688$, above which $F(\epsilon)$ is negative and the angular separation increases with charge, and below which $F(\epsilon)$ is positive and the angular separation decreases with charge. The analysis reveals that the global monopole parameter introduces a non-monotonic behavior in the angular separation with respect to charge. There are critical values of $\epsilon$ ($\epsilon_{0}$ and $\epsilon_{1}$) that determine whether the angular separation increases or decreases with charge. This behavior is distinctly different from the RN black hole case, where the angular separation always increases with charge.
@@
The analysis above focuses on the monotonicity of the angular separation in the small-charge ($\eta\to 0$) and large-charge ($\eta\to 1$) limits. To gain a more comprehensive understanding of its behavior for arbitrary black hole charge, we examine numerically the sign of the partial derivative of the angular separation $s$ with respect to the charge parameter $\eta$ as a function of the global monopole parameter $\epsilon$ and the charge parameter $\eta$, as shown in Fig. 2. Additionally, Figs. 22(b) and 22(c) provide enlarged views of specific regions of Fig. 2(a). The figures clearly reveal that, in addition to the two critical values $\epsilon_{0}$ and $\epsilon_{1}$ found in the analytical investigation, a third critical value $\epsilon_{2}\approx 0.1826576$ emerges from the numerical analysis. These critical points divide the parameter space into four distinct regions that govern the monotonicity and behavior of the angular separation $s$ with respect to the charge parameter ($\eta$): 1. When $\epsilon<\epsilon_{2}$, the partial derivative is always negative, indicating that the angular separation decreases monotonically with charge. 2. When $\epsilon_{2}<\epsilon<\epsilon_{1}$, the angular separation initially decreases, then increases, and subsequently decreases again as the charge increases. This non-monotonic behavior results in the angular separation exhibiting both a minimum and a maximum, as illustrated in Fig. 3. 3. When $\epsilon_{1}<\epsilon<\epsilon_{0}$, the angular separation initially decreases and then increases as the charge increases. This again indicates the existence of a single minimum in the angular separation as a function of charge, as demonstrated in Fig. 3. 4. When $\epsilon>\epsilon_{0}$, the partial derivative is always positive, meaning that the angular separation increases monotonically with charge. These results reveal that the presence of the global monopole introduces more complex behavior in the angular separation as a function of charge. In particular, the angular separation does not always increase with charge, as it does for the RN black hole. Instead, the behavior depends on the global monopole parameter $\epsilon$, with some regions exhibiting a minimum or maximum angular separation as charge varies.
@@
Figure 3: (a) The dependence of angular separation $s$ on the charge parameter $\eta$ when $\epsilon=0.183$, where $\epsilon_{1}<\epsilon<\epsilon_{0}$. (b) The dependence of angular separation $s$ on the charge parameter $\eta$ when $\epsilon=0.182663$, where $\epsilon_{2}<\epsilon<\epsilon_{1}$. In both cases, the angular separation is measured in microarc sec, and $D_{\text{OL}}=10^{10}M$.
III.2.2 The flux ratio
@@
We now numerically analyze the behavior of the flux ratio $\mathcal{R}$ over a broader parameter range. As shown in Fig. 4, the critical value $\eta_{0}=\frac{27}{32}\approx 0.84$ (indicated by the yellow dashed line) divides the parameter space into two distinct regions. When the charge parameter $\eta$ is less than $\eta_{0}$, the flux ratio $\mathcal{R}$ decreases monotonically with the global monopole parameter $\epsilon$. In contrast, when $\eta$ exceeds $\eta_{0}$, the flux ratio $\mathcal{R}$ initially increases and then decreases as $\epsilon$ increases. To illustrate these behaviors explicitly, we show the flux ratio as a function of the global monopole parameter $\epsilon$ for a fixed charge in Fig. 5. From this figure, we observe that when the charge satisfies $\eta\leq\frac{27}{32}$ (i.e., the red and black lines), the flux ratio $\mathcal{R}$ increases as the global monopole parameter $\epsilon$ increases (i.e., as the deficit angle decreases). However, when $\eta>\eta_{0}$ (i.e., the blue line), the flux ratio $\mathcal{R}$ initially increases and then decreases. Additionally, it can also be seen from Fig. 5 that, when the global monopole parameter $\epsilon$ is fixed, the flux ratio $\mathcal{R}$ decreases as the charge (i.e., the parameter $\eta$) increases.

### v3

@@ III Interdependence of gravitational lensing observables on charge and global monopole
Figure 1: Fundamental geometric configuration underlying the phenomenon of gravitational lensing. $b$ is the impact parameter, $r_{0}$ represents the minimum distance between the light ray and the lens, and $I$ and $L$ denote the positions of the image and the lens. $S$ and $O$ represent the source and observer.
The image formed by photons that encircle a lens $n$ times is referred to as the $n$th relativistic image. Using the lens equation (3.1) and the analytical expression for the deflection angle (2.23), the angular position of the $n$th relativistic image in the strong field regime is given by Bozza (2002)
$\theta_{n}=\theta^{0}_{n}+\frac{b_{m}e_{n}(\beta-\theta^{0}_{n})D_{\text{OS}}}{\bar{a}D_{\text{LS}}D_{\text{OL}}},$ (3.2)
where
$\theta^{0}_{n}=\frac{b_{m}}{D_{\text{OL}}}(1+e_{n}),$ (3.3)
and
$e_{n}=e^{\frac{\bar{b}-2n\pi}{\bar{a}}}.$ (3.4)
The magnification of the $n$th relativistic image is then given by Bozza (2002)
$\mu_{n}=\left(\frac{\beta}{\theta}\frac{d\beta}{d\theta}\right)^{-1}\bigg{|}_{\theta_{n}^{0}}=e_{n}\frac{b_{m}^{2}\left(1+e_{n}\right)D_{\text{OS}}}{\bar{a}\beta D_{\text{OL}}^{2}D_{\text{LS}}},$ (3.5)
When the source, lens, and observer are nearly perfectly aligned, i.e., $\beta=0$, the angular radius of the $n$th Einstein ring is given by Bozza (2002)
$\theta_{n}^{E}=\left(1-\frac{D_{\text{OS}}}{D_{\text{LS}}D_{\text{OL}}}\frac{b_{m}e_{n}}{\bar{a}}\right)\theta^{0}_{n}.$ (3.6)
Since $b_{m}\ll D_{\text{OL}}$, the size of the Einstein ring can be well approximated by $\theta^{0}_{n}$ as given in Eq. (3.3).
In addition to the observables discussed above, there are three other essential observables in gravitational lensing.
First, there is the angular separation. It is hypothesized that the outermost (first) image in the series of images stands apart from the others, being uniquely discernible, while the remaining images nearly overlap and are difficult to distinguish. Therefore, we introduce an observable quantity: the angular separation, denoted as $s$, which quantifies the difference between the angular position of the first image and that of the subsequent ones. This quantity is defined as Bozza (2002)
$s=\theta_{1}-\theta_{\infty}\approx\frac{b_{m}}{D_{\text{OL}}}e_{1}.$ (3.7)
Second, there is the flux ratio. The flux ratio, also known as the relative magnification, quantifies the brightness contrast between the primary image and the cumulative flux of all subsequent relativistic images. It is given by Bozza (2002)
@@
Third, there is the time delay. In the strong field regime, photons can orbit a black hole multiple times, leading to the formation of multiple relativistic images. Due to the differences in the path lengths taken by these photons, the corresponding light travel times are also different. The time delay between the $n$th and $k$th relativistic images, which are formed by photons traveling along different paths on the same side of the lensing object, is given by Bozza and Mancini (2004)
$\Delta T_{n,k}^{s}=b_{m}\left[2\pi\left(n-k\right)+2\sqrt{2}\left(\sqrt{e_{k}}-\sqrt{e_{n}}\right)\pm\frac{\sqrt{2}D_{\text{OS}}\beta}{\bar{a}D_{\text{LS}}}\left(\sqrt{e_{k}}-\sqrt{e_{n}}\right)\right],$ (3.11)
where the positive (negative) sign applies when both images are located on the same (opposite) side of the source. For images on opposite sides of the lens, the time delay is given by Bozza and Mancini (2004)
$\Delta T_{n,k}^{o}=b_{m}\left[2\pi\left(n-k\right)+2\sqrt{2}\left(\sqrt{e_{k}}-\sqrt{e_{n}}\right)+\frac{\sqrt{2}D_{\text{OS}}\beta}{\bar{a}D_{\text{LS}}}\left(\sqrt{e_{k}}-\sqrt{e_{n}}\right)-\frac{2D_{\text{OS}}\beta}{D_{\text{LS}}}\right].$ (3.12)
@@
By leveraging the coefficients $\bar{a}$ and $\bar{b}$ of the deflection angle, we now derive explicit analytical expressions for the key observables in gravitational lensing, including the angular position and the magnification of the $n$th relativistic image, the angular radius of the Einstein ring, the angular separation, the flux ratio, and the time delay. Previous studies have primarily relied on numerical approaches to analyze these observables in specific scenarios. For instance: Reference Bozza (2002) conducted numerical computations for the angular separation $s$ and the flux ratio $\mathcal{R}$ of an RN black hole. Reference Cheng and Man (2011) focused on a Schwarzschild black hole with a global monopole, again without explicit analytical expressions for the observables. In this work, we go beyond previous studies by deriving explicit analytical expressions for both $s$ and $\mathcal{R}$ for an RN black hole with a global monopole, facilitating a direct evaluation of how these observables respond to variations in charge ($\eta$) and the global monopole parameter ($\epsilon$). To achieve a comprehensive understanding, we conduct both analytical and numerical analyses, examining both the individual effects of the global monopole and charge on the observables, as well as the combined impact of charge and the global monopole, revealing their interplay in modifying strong field lensing behavior.
III.1 Angular separation
@@
Secondly, we investigate the dependence of the angular separation on the charge for a fixed global monopole parameter. First, when there is charge but no global monopole (i.e., $\epsilon=1$), the angular separation is given by the expression
@@
where $w$ represents the dimensionless radius of the photon sphere, $\bar{r}_{m}(1,\eta)$. Furthermore, as shown in Appendix A, the derivative of the angular separation with respect to $\eta$ is always positive. This indicates that the angular separation monotonically increases with increasing charge. Now, we analyze the behavior of the angular separation $s$ in the presence of both charge and a global monopole. Directly analyzing Eq. (3.13) is challenging due to its complexity. To understand the effects of both charge and the global monopole, we expand the angular separation expression for small and large charge values, respectively. For small charge ($\eta\ll 1$), Eq. (3.13) can be expanded as
@@
The dependence of the angular separation on charge is determined by the coefficient of the $\eta$-dependent term. There exists a critical value of the global monopole parameter, given by $\epsilon_{0}=\left[\frac{51\sqrt{3}-88}{2\pi\left(7\sqrt{3}-12\right)}\right]^{2}\approx 0.1833742$, which separates two behaviors: For $\epsilon>\epsilon_{0}$, coefficient of the $\eta$-dependent term is positive, and the angular separation increases with charge. Conversely, for $\epsilon<\epsilon_{0}$, the coefficient is negative, and the angular separation decreases with charge. For large charge ($\eta\to 1$), the angular separation (Eq. (3.13)) behaves as
@@
where $H(\epsilon)$ and $F(\epsilon)$ are functions of the global monopole parameter $\epsilon$. Due to the complexity of these expressions, their explicit forms are provided in Appendix B. We find that, similar to the small-charge case, there exists a critical value $\epsilon_{1}\approx 0.1826688$, above which $F(\epsilon)$ is negative and the angular separation increases with charge, and below which $F(\epsilon)$ is positive and the angular separation decreases with charge. The analysis reveals that the global monopole parameter introduces a nonmonotonic behavior in the angular separation with respect to charge. There are critical values of $\epsilon$ ($\epsilon_{0}$ and $\epsilon_{1}$) that determine whether the angular separation increases or decreases with charge. This behavior is distinctly different from the RN black hole case, where the angular separation always increases with charge.
@@
The analysis above focuses on the monotonicity of the angular separation in the small-charge ($\eta\to 0$) and large-charge ($\eta\to 1$) limits. To gain a more comprehensive understanding of its behavior for arbitrary black hole charge, we examine numerically the sign of the partial derivative of the angular separation $s$ with respect to the charge parameter $\eta$ as a function of the global monopole parameter $\epsilon$ and the charge parameter $\eta$, as shown in Fig. 2. Additionally, Figs. 22(b) and 22(c) provide enlarged views of specific regions of Fig. 2(a). The figures clearly reveal that, in addition to the two critical values $\epsilon_{0}$ and $\epsilon_{1}$ found in the analytical investigation, a third critical value $\epsilon_{2}\approx 0.1826576$ emerges from the numerical analysis. These critical points divide the parameter space into four distinct regions that govern the monotonicity and behavior of the angular separation $s$ with respect to the charge parameter ($\eta$):
  1. 1.
    When $\epsilon<\epsilon_{2}$, the partial derivative is always negative, indicating that the angular separation decreases monotonically with charge.
  2. 2.
    When $\epsilon_{2}<\epsilon<\epsilon_{1}$, the angular separation initially decreases, then increases, and subsequently decreases again as the charge increases. This nonmonotonic behavior results in the angular separation exhibiting both a minimum and a maximum, as illustrated in Fig. 3.
  3. 3.
    When $\epsilon_{1}<\epsilon<\epsilon_{0}$, the angular separation initially decreases and then increases as the charge increases. This again indicates the existence of a single minimum in the angular separation as a function of charge, as demonstrated in Fig. 3.
  4. 4.
    When $\epsilon>\epsilon_{0}$, the partial derivative is always positive, meaning that the angular separation increases monotonically with charge.
These results reveal that the presence of the global monopole introduces more complex behavior in the angular separation as a function of charge. In particular, the angular separation does not always increase with charge, as it does for the RN black hole. Instead, the behavior depends on the global monopole parameter $\epsilon$, with some regions exhibiting a minimum or maximum angular separation as charge varies.
@@
Figure 3: (a) The dependence of angular separation $s$ on the charge parameter $\eta$ when $\epsilon=0.183$, where $\epsilon_{1}<\epsilon<\epsilon_{0}$. (b) The dependence of angular separation $s$ on the charge parameter $\eta$ when $\epsilon=0.182663$, where $\epsilon_{2}<\epsilon<\epsilon_{1}$. In both cases, the angular separation is measured in microarcsecond, and $D_{\text{OL}}=10^{10}M$.
III.2 Flux ratio
@@
We now numerically analyze the behavior of the flux ratio $\mathcal{R}$ over a broader parameter range. As shown in Fig. 4, the critical value $\eta_{0}=\frac{27}{32}\approx 0.84$ (indicated by the yellow dashed line) divides the parameter space into two distinct regions. When the charge parameter $\eta$ is less than $\eta_{0}$, the flux ratio $\mathcal{R}$ decreases monotonically with the global monopole parameter $\epsilon$. In contrast, when $\eta$ exceeds $\eta_{0}$, the flux ratio $\mathcal{R}$ initially increases and then decreases as $\epsilon$ increases. To illustrate these behaviors explicitly, we show the flux ratio as a function of the global monopole parameter $\epsilon$ for a fixed charge in Fig. 5. From this figure, we observe that when the charge satisfies $\eta\leq\frac{27}{32}$ (i.e., the red and black lines), the flux ratio $\mathcal{R}$ increases as the global monopole parameter $\epsilon$ increases (i.e., as the deficit angle decreases). However, when $\eta>\eta_{0}$ (i.e., the blue line), the flux ratio $\mathcal{R}$ initially increases and then decreases. Furthermore, it can also be seen from Fig. 5 that, when the global monopole parameter $\epsilon$ is fixed, the flux ratio $\mathcal{R}$ decreases as the charge (i.e., the parameter $\eta$) increases.
@@
III.3 Magnification of the first image
The magnification of the first image, denoted as $\mu_{1}$ can be obtained from Eq. (3.5) as follows
$\mu_{1}(\epsilon,\eta)=e_{1}\frac{b_{m}^{2}\left(1+e_{1}\right)D_{\text{OS}}}{\bar{a}\beta D_{\text{OL}}^{2}D_{\text{LS}}},$ (3.23)
where $\bar{a}$ and $b_{m}$ represent $\bar{a}(\epsilon,\eta)$ and $b_{m}(\epsilon,\eta)$, respectively, and $e_{1}=e^{\frac{\bar{b}-2\pi}{\bar{a}}}$.
Direct analytical examination of the behavior of $\mu_{1}$ is challenging, so we perform a numerical analysis. The sign of the first-order partial derivative of $\mu_{1}$ with respect to the charge parameter $\eta$ is presented in Fig. 6, with Fig. 6(b) providing an enlarged view of specific regions from Fig. 6(a). These figures reveal the existence of two critical values $\epsilon_{3}\approx 0.4674$ and $\epsilon_{4}\approx 0.4295$, which divide the parameter space into three distinct regions, each exhibiting different monotonic behaviors of the magnification $\mu_{1}$ with respect to the charge parameter $\eta$.
  1. 1.
    When $\epsilon<\epsilon_{4}$, the partial derivative of the magnification with respect to charge is always negative, indicating that the magnification decreases monotonically as the charge increases.
  2. 2.
    When $\epsilon_{4}<\epsilon<\epsilon_{3}$, the magnification initially decreases and then increases as the charge increases, indicating the existence of a single minimum in the magnification as a function of charge, as demonstrated in Fig. 7.
  3. 3.
    When $\epsilon>\epsilon_{3}$, the partial derivative is always positive, meaning that the magnification increases monotonically with charge.
Additionally, direct numerical calculations show that the partial derivative of $\mu_{1}$ with respect to the global monopole parameter $\epsilon$ is always negative, suggesting that the magnification decreases as $\epsilon$ increases.
Figure 6: (a) The sign of the first-order partial derivative of the magnification $\mu_{1}$ of the first image with respect to the charge parameter $\eta$, where positive values are shown in red and negative values in blue. (b) An enlarged view of Fig. 6(a) in the region $\epsilon\in\left[0.42,0.48\right]$. The white and green dashed lines correspond to the critical values $\epsilon_{3}\approx 0.4674$ and $\epsilon_{4}\approx 0.4295$, respectively.
Figure 7: The magnification $\mu_{1}$ of the first image as a function of the charge parameter $\eta$ for a fixed global monopole parameter $\epsilon$. The actual magnification is given by $\mu_{1}\times 10^{-20}$. Here $\epsilon=0.45$, and $D_{\text{OL}}=D_{\text{LS}}=10^{10}M$.
III.4 Size of the first Einstein ring
The angular radius of the first Einstein ring can be well approximated by $\theta^{0}_{1}$, as given in Eq. (3.3). Substituting the gravitational lensing coefficients $\bar{a},\bar{b}$ from Eqs. (2.26) and (2.27), along with the critical impact parameter $b_{m}$ from Eq. (2.10), into Eq. (3.3), we obtain the explicit expression
$\theta_{1}^{E}\left(\epsilon,\eta\right)=\frac{M\bar{r}_{m}^{2}}{D_{OL}\sqrt{\epsilon\bar{r}_{m}^{2}-2\bar{r}_{m}+\eta}}\left\{1+\left[\frac{8\left(3\bar{r}_{m}-4\eta\right)^{3}}{\bar{r}_{m}^{2}\left(\bar{r}_{m}-\eta\right)^{2}}\left(2\sqrt{\bar{r}_{m}-\eta}-\sqrt{3\bar{r}_{m}-4\eta}\right)^{2}\right]\exp\left(\frac{-3\pi\sqrt{\epsilon\bar{r}_{m}^{2}-2\eta}}{\bar{r}_{m}}\right)\right\},$ (3.24)
where $\bar{r}_{m}$ is $\bar{r}_{m}(\epsilon,\eta)$. Due to the complexity of this expression, analytical analysis is intractable; therefore, we perform a numerical investigation. Using the same method as before, we compute the partial derivatives of $\theta_{1}^{E}$ with respect to $\epsilon$ and $\eta$ and find both to be negative. This indicates that the radius of the first Einstein ring $\theta_{1}^{E}$ always increases as the deficit solid angle $\Delta$ due to the global monopole grows (i.e., as $\epsilon$ decreases), irrespective of the charge of the black hole. Conversely, $\theta_{1}^{E}$ always decreases with increasing charge, regardless of the presence of a global monopole. To visualize these behaviors, Fig. 8 plots $\theta_{1}^{E}$ as a function of the black hole charge parameter $\eta$ for several fixed values of $\epsilon$. Each curve shows a monotonic decrease in $\theta_{1}^{E}$ with increasing $\eta$. Additionally, comparisons across curves reveal that a larger deficit angle (i.e., smaller $\epsilon$) leads to a larger Einstein ring.
Figure 8: Angular radius of the first Einstein ring $\theta_{1}^{E}$ as a function of the charge parameter $\eta$, for various values of the global monopole parameter $\epsilon$. The red, black, purple, and blue curves correspond to $\epsilon=1$, $0.9$, $0.8$, and $0.7$, respectively. The angular radius is measured in microarcseconds. The lens-observer distance is set to $D_{\text{OL}}=10^{10}M$.
III.5 Time delay between the first and second images
According to Eq. (3.11), the time delay between the first and second relativistic images, both formed on the same side of the lens and the source by photons traversing different paths, is given by
$\Delta T_{2,1}^{s+}(\epsilon,\eta)=\frac{\bar{r}_{m}^{2}}{\sqrt{\epsilon\bar{r}_{m}^{2}-2\bar{r}_{m}+\eta}}\left[2\pi+2\sqrt{2}\left(\sqrt{e_{1}}-\sqrt{e_{2}}\right)+\frac{\sqrt{2}D_{\text{OS}}\beta}{\bar{a}D_{\text{LS}}}\left(\sqrt{e_{1}}-\sqrt{e_{2}}\right)\right].$ (3.25)
Here, $\bar{r}_{m}=\bar{r}_{m}(\epsilon,\eta)$, $\bar{a}=\bar{a}(\epsilon,\eta)$, and $e_{n}$ is defined in Eq. (3.4). Due to the complexity of this expression, we carry out a numerical investigation. Following the same methodology as in earlier sections, we compute the partial derivatives of $\Delta T_{2,1}^{s+}$ with respect to $\epsilon$ and $\eta$. Our analysis reveals that both derivatives are consistently negative. This implies that the time delay increases with the deficit angle $\Delta$ due to the global monopole (i.e., as $\epsilon$ decreases), regardless of whether the black hole is charged. Conversely, $\Delta T_{2,1}^{s+}$ decreases monotonically as the charge parameter $\eta$ increases, independently of the value of $\epsilon$. To visualize these dependencies, Fig. 9 displays $\Delta T_{2,1}^{s+}$ as a function of the charge parameter $\eta$ for various fixed values of the global monopole parameter $\epsilon$. Each curve clearly shows a decrease in time delay with increasing charge, while comparison across curves indicates that larger deficit angles (smaller $\epsilon$) yield longer time delays.
Figure 9: Dimensionless time delay $\Delta T_{2,1}^{s+}/M$ between the first and second relativistic images as a function of the charge parameter $\eta$, for several fixed values of the global monopole parameter $\epsilon$. The solid red, black, purple, and blue lines correspond to $\epsilon=1$, $0.9$, $0.8$, and $0.7$, respectively. The lens-observer distance is set to $D_{\text{OL}}=10^{10}M$, with $D_{\text{LS}}=D_{\text{OL}}=D_{\text{OS}}/2$.